Determine whether each statement is sometimes, always, or never true.
a. √(a2+b2) is a real number.

Answer by stanbon(75887)

Determine whether each statement is sometimes, always, or never true.
a. √(a2+b2) is a real number.
----------
a^2 is always >= zero
b^2 is always >= zero
--------
Therefore a^2+b^2 >=0
Therefore sqrt(a^2+b^2) is always a Real Number.
================
Cheers,
Stan H.